# SoK: Root Cause of $1 Billion Loss in Smart Contract Real-World Attacks via a Systematic Literature Review of Vulnerabilities

### v1

@@ Abstract
The Ethereum ecosystem, despite its maturity, continues to witness catastrophic attacks, with billions of dollars in assets lost annually. In response, a significant body of research has focused on identifying and mitigating smart contract vulnerabilities. However, these efforts predominantly focus on implementation-level bugs, leaving a critical gap between academic understanding of vulnerabilities and the root causes of real-world high-impact financial losses. We employ a two-pronged methodology: first, a systematic literature review of 71 academic papers to build a comprehensive and up-to-date catalog of 24 active and 5 deprecated vulnerabilities as understood by the research community. Second, we conduct an in-depth, empirical analysis of 50 of the most severe real-world exploits between 2022 and 2025, collectively incurring over $1.09B in losses, to identify their true root causes. We introduce the concept of “exploit chains” by revealing that many incidents are not caused by isolated vulnerabilities but by combinations of human, operational, and economic design flaws that link with implementation bugs to enable an attack. Our analysis yields insights on how DApps are exploited in practice, leading to a novel, four-tier root-cause framework that moves beyond code-level vulnerabilities. We find that real-world successful attacks on Ethereum (and related networks) trace back to one of the four tiers of (1) protocol logic design, (2) lifecycle and governance, (3) external dependencies, and (4) traditional implementation bugs (classic smart contract vulnerabilities). We investigate the suitability of this multi-tier incident root-cause framework via a case study.
@@ I Introduction
Table I presents a review of three Systematization of Knowledge (SoK) and ten Systematic Literature Review (SLR) papers on Decentralized Application (DApp) security. Surveys and SoKs classify tens of vulnerabilities across different layers of the blockchain stack (e.g., network, smart contract, and protocol layers) [12, 17, 18]. These taxonomies have become increasingly granular in classifying weaknesses [17] and mapping them to formal verification properties [15], or aligning them with frameworks such as NIST bugs [14]. However, our review unveils a few gaps in this literature: First, prevention gap: only 4 of 13 works (S4, S9, S12, and S13) address prevention methods, most of which focus exclusively on detection. Second, the validation gap: merely 3 works validate their findings against real-world exploits—S3, S9 and S12 or analyze case studies, and exploits, while others rely on synthetic datasets. Third, the taxonomy evolution shows a shift from simple categorization to sophisticated multi-layer frameworks (4L, 5L), yet none adopt a root-cause approach based on real incidents.
Real-World Incidents Tell-Tale
At 08:56 UTC on 13 March 2023, the Euler Finance lending market was solvent; four minutes later, it was $ million in deficit [24, 25]. Nothing in its Solidity source code could be attributed to common smart contract vulnerabilities, no unchecked casts, no arithmetic overflow, or no reentrancy bug. Instead, an unguarded auxiliary function, a liquidation discount, and a flash loan combined into an unexpected profit engine to exploit the contract. Several months earlier, the Nomad bridge was emptied of $M because a single initializer function had been called with the wrong argument during an upgrade, resetting its trusted root to zero address and turning every message into valid [26]. Episodes like these are not rare accidents [5]. The vast majority of existing literature and automated tooling remains heavily focused on implementation-level weaknesses, the “known unknowns” [12, 17, 15]. Yet, as incidents like the $M Euler Finance exploit [24] show, financially devastating attacks may not merely stem from coding mistakes. In the mentioned protocol’s case, the attack succeeded as the liquidation mechanism was flawed. The code executed exactly as intended; it was the intention itself that created the vulnerability. This incident is not an isolated case. It highlights a systemic issue: while literature focuses on finding bugs in the code, a large-scale analysis of real-world attacks and comparing the exploited vulnerabilities in these incidents and their alignment with the literature is required.
@@
RQ2: What factors, other than mere coding vulnerability patterns (identified in RQ1), are evident in the incidents and contribute to the success of the attack?
@@ II-A Academic Literature Study Protocol
To answer RQ1, we conducted a systematic literature review (SLR) to obtain a comprehensive and structured understanding of Ethereum smart contract vulnerabilities identified in the academic literature. Following Kitchenham’s guidelines [27], our review process includes study identification (Section II-A1) and study selection (Section II-A2).
@@ II-A2 Study Selection
Following the initial identification of papers, we removed duplicates and filtered them based on the exclusion criteria in Table III. Specifically, language (EC2) and publication year 2018–2024 (EC1) filtering reduced the count to papers. Title and abstract screening using criteria EC3 (Ethereum focus), EC5, and EC6 (topic relevance) yielded papers. Full-text assessment based on availability (EC4) further refined the selection to papers. Finally, quality-based evaluation via manual expert empirical validation and identification of novel contributions resulted in papers for in-depth analysis. Figure 1 illustrates this process.
@@ III-A6 DoS Attack via Owner Account
Contracts should 1) restrict access to critical functions, 2) use multi-signature approvals for transactions, and 3) avoid reliance on a single account for key operations [59].
@@ III-A9 Insecure Randomness
Smart contracts often depend on external facts such as token prices, weather data, sports scores to make irreversible decisions. An oracle is the component that transports those facts on chain (or acts as an service contract to provide data from other on-chain services). Recent studies show that integrity of the data can be manipulated during different stages such as creation, forwarding, election, consumption during its lifecycle [4]. We identified two settings where data feeds are manipulatable:
@@
Centralised off‑chain oracles. System model: one or a few trusted reporters sign values off-chain and publish them on-chain. Attacker model: compromise, bribe, or simply delay any reporter to feed stale or false data [80].
@@
Robust temporal aggregation. Use multi‑block medians or long‑window time-weighted average pricing (TWAP) taken to smooth out single‑block distortions [4].
@@
On‑chain invariants and circuit breakers. Simple require or assert statements that cap per‑block price changes, or tool‑generated invariants such as those from Trace2Inv [81] can stop attacks at runtime.
@@
This vulnerability occurs when the contract’s logic depends on the value of block.timestamp (the block’s timestamp). This vulnerability arises due to the clock drift of distributed systems, which allows miners to adjust the timestamp within a specific range, typically seconds in Ethereum [118]. Smart contract condition checks may be impacted by this manipulation, even if they have nothing to do with producing random numbers. For instance, consensus participants can alter the end time of events like bids or auctions to their advantage [119]. Validators can reduce or lengthen waiting periods or lock-up periods before specific activities are initiated in a contract, changing the outcome to the advantage of a particular group [43]. Additionally, it is a typical manipulation for miners to reorganize and modify the timing of transactions, particularly in contracts that depend on the precise timing or order of transactions. Changing the timestamp may also alter the outcome of computations that depend on time, causing the contract to behave inconsistently. Miners may modify the voting schedule and outcomes to benefit themselves or a certain group, which could have an effect on governance ecosystems [120].
@@ III-B1 Upgradable Contracts
Smart contracts often incorporate upgradability mechanisms through registries, logic contracts, and proxies to manage immutability constraints of the code section of the contract [221]. Although early analyses classified upgradability as a potential vulnerability [12], recent advancements and governance mechanisms have demonstrated that upgradability is better understood as an inherent technical challenge rather than a security flaw. On-chain governance proposals [222, 223] now enable decentralized control over upgrades, reducing the risk of unilateral malicious actions. Moreover, research efforts such as SoliNomic [224] have introduced formal refinement-based approaches to structuring the evolution of decentralized organizations (for correct and consistent upgrades). Therefore, upgradability, when properly designed, enhances the adaptability of decentralized systems without compromising security.
@@ III-B5 Call‑Stack Depth Limit
To answer RQ2, we conducted a comprehensive analysis of significant smart contract exploits that occurred between and early . The collection, selection, and analysis protocol for these attacks is presented in Section II-B. Table V presents the main attributes of these attacks, including date, volume of financial loss, ecosystem, application type, and the exploited vulnerabilities (identified through manual expert analysis and based on active vulnerabilities in Section III-A). We present more information of each attack and its attributes (victim contact and attacker addresses, etc.) along with a summary of the technical details of each attack in our accompanying repository333https://github.com/HadisRe/SoK-Root-Cause-of-Smart-Contract-Incidents/.
Our insights reveal that while real-world successful attacks involve technical weaknesses that could be attributed to classes of vulnerabilities identified by the literature, there are dominant patterns in the incidents that have been unnoticed by the literature. Access control vulnerabilities and price manipulation emerged as the most frequent and financially devastating attack vectors, accounting for a combined loss of $ million in our dataset (see Table VI). Access control-related incidents involve exploit methods beyond mere coding flaws (e.g., lack or misplacement of require(msg.sender==owner) checks) and involve human-centric factors as well (see Insight IV). The high frequency of price manipulation attacks underscores the critical importance of secure, manipulation-resistant oracles (Insight IV). We also observed that while classic vulnerabilities like reentrancy persist, attackers are evolving their techniques, employing sophisticated variants like read-only reentrancy and targeting new EVM features such as transient storage (Insight IV), opening up novel exploit avenues.
@@
For a detailed review report of each incident, refer to our accompanying repository444https://github.com/HadisRe/SoK-Root-Cause-of-Smart-Contract-Incidents/ providing in-depth analysis.
@@ V A Multi-Tier Root-Cause Framework for Smart Contract Exploits
| Operational gates | Multi‑sig + two‑step pause upgrade un‑pause. | Human review spots nonsense state. |
@@ Conclusion
This study unites the literature’s focus on implementation bugs with the practitioner’s day‑to‑day experience of successful attacks on Ethereum‑based protocols. Through a systematic literature review of peer‑reviewed papers and dissecting fifty publicly disclosed exploits that caused a total loss of $ between 2022 and 2025, we show that less than one‑third of recorded losses can be traced to an isolated coding mistake. The remainder cascaded from flawed incentive design, clumsy governance decisions, unsafe assumptions about external dependencies, or exploit chains-factors that mainstream static and dynamic analyzers rarely take into account. To capture this, we introduced a four‑tier root‑cause framework ranging from economic logic at tier 1 to programming mistakes at tier 4; together with our labeled incident dataset and mitigation catalog, the framework equips auditors, tool builders and protocol designers with a map of the modern threat surface based on real incidents.

### v2

# SoK: Root Causes of $1 Billion Loss in Smart Contract Real-World Attacks via a Systematic Literature Review of Vulnerabilities

@@ Abstract
While catastrophic attacks on Ethereum persist, vulnerability research remains fixated on implementation-level smart contract bugs, creating a gap between academic understanding of vulnerabilities and the true root causes of high-impact, real-world incidents. To address this, we employ a two-pronged methodology: first, a systematic literature review of 71 academic papers to build a catalog of 24 active and 5 deprecated vulnerabilities. Second, we conduct an in-depth, empirical analysis of 50 of the most severe real-world attacks between 2022 and 2025, collectively incurring over $1.09B in losses, to identify their root causes. We introduce the concept of “exploit chains” by revealing that many incidents are not caused by isolated vulnerabilities but by combinations of human, operational, and economic design flaws that link with implementation bugs to enable an attack. Our analysis yields insights on how decentralized applications are exploited in practice, leading to a novel, four-tier root-cause framework that moves beyond code-level vulnerabilities. We find that real-world successful attacks on Ethereum (and related networks) trace back to one of the four tiers of (1) protocol logic design, (2) lifecycle and governance, (3) external dependencies, and (4) classic smart contract vulnerabilities. We investigate the suitability of this multi-tier incident root-cause framework via a case study.
@@ I Introduction
Table I presents a review of three Systematization of Knowledge (SoK) and ten Systematic Literature Review (SLR) papers on Decentralized Application (DApp) security. Surveys and SoKs classify tens of vulnerabilities across different layers of the blockchain stack (e.g., network, smart contract, and protocol layers) [12, 17, 18]. These taxonomies have become increasingly granular in classifying weaknesses [17] and mapping them to formal verification properties [15], or aligning them with frameworks such as NIST bugs [14]. However, our review unveils a few gaps in this literature: First, prevention gap: only 4 of 13 works (S4, S9, S12, and S13) address prevention methods, most of which focus exclusively on detection. Second, the validation gap: merely 3 works validate their findings against real-world exploits—S3 (case studies), S9 (181 exploits) and S12 (92 exploits), while others rely on synthetic datasets. Third, the taxonomy evolution shows a shift from simple categorization to sophisticated multi-layer frameworks (4L, 5L), yet none adopt a root-cause approach based on real incidents.
Telltale Signs from Real-World Incidents
At 08:56 UTC on 13 March 2023, the Euler Finance lending market was solvent; four minutes later, it was $ million in deficit [24, 25]. Nothing in its Solidity source code could be attributed to common smart contract vulnerabilities, no unchecked casts, no arithmetic overflow, or no reentrancy bug. Instead, an unguarded auxiliary function, a liquidation discount, and a flash loan combined into an unexpected profit engine to exploit the contract. Several months earlier, the Nomad bridge was emptied of $M because a single initializer function had been called with the wrong argument during an upgrade, resetting its trusted root to zero address and turning every message into valid [26]. Episodes like these are not rare accidents [5]. The vast majority of existing literature and automated tooling remains heavily focused on implementation-level weaknesses, the “known unknowns” [12, 17, 15]. Yet, as incidents like the $M Euler Finance exploit [24] show, financially devastating attacks may not merely stem from coding mistakes. In the mentioned protocol’s case, the attack succeeded as the liquidation mechanism was flawed. The code executed exactly as intended; it was the intention itself that created the vulnerability. This incident is not an isolated case. It highlights a systemic issue: while literature focuses on finding bugs in the code, a large-scale analysis of real-world attacks and comparing the exploited vulnerabilities in these incidents and their alignment with the literature is required. Thus, we propose the following research questions:
@@
RQ2: What factors, other than mere coding vulnerability patterns (identified in RQ1), are evident in the real-world incidents and contribute to the success of the attacks?
@@ II-A Academic Literature Study Protocol
To answer RQ1, we conducted a systematic literature review (SLR) to obtain a comprehensive and structured understanding of Ethereum smart contract vulnerabilities from academic literature. Following Kitchenham‘’s guidelines [27], our review process includes study identification (Section II-A1) and study selection (Section II-A2).
@@ II-A2 Study Selection
Following the initial identification of papers, we removed duplicates and filtered them based on the exclusion criteria in Table III. Specifically, language (EC2) and publication year 2018–2024 (EC1) filtering reduced the count to papers. Title and abstract screening using criteria EC3 (Ethereum focus), EC5, and EC6 (topic relevance) yielded papers. Full-text assessment based on availability (EC4) further refined the selection to papers. Finally, quality-based evaluation via manual review of the paper and identification of novel contributions resulted in papers for in-depth analysis. Figure 1 illustrates this process.
@@ III-A6 DoS Attack via Owner Account
Contracts should restrict access to critical functions multi-signature approvals for transactions to avoid reliance on a single account for key operations [59].
@@ III-A9 Insecure Randomness
Smart contracts often depend on external facts such as token prices, weather data, and sports scores to make irreversible decisions. An oracle is the component that transports those facts on chain (or acts as an service contract to provide data from other on-chain services). Recent studies show that integrity of the data can be manipulated at stages of its lifecycle [4]. We identified two settings where data feeds of these oracles are manipulable:
@@
Off‑chain oracles. System model: trusted reporters sign values off-chain and the oracle publishes them on-chain. Attacker model: compromise, bribe, or simply delay reporters to feed stale or false data [80].
@@
Temporal aggregation. To deter price manipulation in DEX-based oracles, use multi‑block medians or long‑window time-weighted average pricing (TWAP) taken to smooth out single‑block distortions [4].
@@
On‑chain invariants and circuit breakers. Use invariants (in form of require statements) that cap per‑block price changes to stop attacks at runtime [81] .
@@
This vulnerability occurs when the contract’s logic depends on the value of block.timestamp. Due to the clock drift of distributed systems, miners are allowed to adjust the timestamp within a specific range, typically seconds in Ethereum [118]. Smart contract condition checks may be impacted by this manipulation, even if they have nothing to do with producing random numbers. For instance, consensus participants can alter the end time of events like bids or auctions [119]. Validators can lengthen waiting or lock-up periods before specific activities to change the outcome to the advantage of a particular group [43]. Additionally, it is a typical manipulation for miners to reorganize and modify the timing of transactions, particularly in contracts that depend on the precise timing or order of transactions. Changing the timestamp may also alter the outcome of computations that depend on time, causing the contract to behave inconsistently. Miners may modify the voting schedule and outcomes to benefit themselves or a certain group, which could have an effect on governance ecosystems [120].
@@ III-B1 Upgradable Contracts
Smart contracts often incorporate upgradability mechanisms through logic contracts and proxies to manage immutability constraints of the code section of the contract [221]. Although early analyses classified upgradability as a potential vulnerability [12], recen advancements in upgradability patterns and governance mechanisms have demonstrated that this is better understood as a technical challenge rather than a security weakness. On-chain governance proposals [222] now enable decentralized control over upgrades, reducing the risk of unilateral malicious actions. Moreover, research efforts such as SoliNomic [223] have introduced formal refinement-based approaches to structuring the evolution of decentralized organizations (for correct and consistent upgrades). Therefore, upgradability, when properly designed, enhances the adaptability of DApps without compromising security.
@@ III-B5 Call‑Stack Depth Limit
To answer RQ2, we conducted a comprehensive analysis of significant smart contract exploits that occurred between and early . The collection, selection, and analysis protocol for these attacks is presented in Section II-B. Table V presents the main attributes of these attacks, including date, volume of financial loss, ecosystem, application type, and the exploited vulnerabilities (identified through manual expert analysis and based on active vulnerabilities in Section III-A). We present more information of each attack and its attributes (victim contact and attacker addresses, etc.) along with a summary of the technical details of each attack in our accompanying repository.
Our insights reveal that while real-world successful attacks involve technical weaknesses attributed to classes of known vulnerabilities (Section III-A), there are dominant patterns in the incidents that have been unnoticed by the literature. Access control vulnerabilities and price manipulation emerged as the most frequent and financially devastating attack vectors, accounting for a combined loss of $ million in our dataset (see Table VI). Access control-related incidents involve exploit methods beyond mere coding flaws (e.g., lack or misplacement of require(msg.sender==owner) checks) and include human-centric factors as well (see Insight IV). The high frequency of price manipulation attacks underscores the critical importance of secure, manipulation-resistant oracles (Insight IV). We also observed that while classic vulnerabilities like reentrancy persist, attackers are employing more sophisticated variants like read-only reentrancy and targeting new EVM features such as transient storage (Insight IV), opening up novel exploit avenues.
@@
For a detailed review of each incident, refer to our accompanying repository333https://github.com/HadisRe/SoK-Root-Cause-of-Smart-Contract-Incidents providing in-depth analysis.
@@ V A Multi-Tier Root-Cause Framework for Smart Contract Exploits
| Operational gates | Multi‑sig + two‑step pause upgrade un‑pause. | Human review spots violating states. |
@@ Conclusion
This study unites the literature’s focus on implementation bugs with the practitioner’s day‑to‑day experience of successful attacks on Ethereum‑based protocols. Through a systematic literature review of peer‑reviewed papers and dissecting publicly disclosed exploits that caused a total loss of $ between 2022 and 2025, we show that less than one‑third of recorded losses can be traced to an isolated coding mistake. The remainder cascaded from flawed incentive design, clumsy governance decisions, unsafe assumptions about external dependencies, or exploit chains-factors that mainstream static and dynamic analyzers rarely take into account. To capture this, we introduced a four‑tier root‑cause framework ranging from economic logic at tier 1 to programming mistakes at tier 4; together with our labeled incident dataset and mitigation catalog, the framework equips auditors, tool builders and protocol designers with a map of the modern threat surface based on real incidents.